OpenStudy (anonymous):

Solve the square root of 12 plus 4 x is greater than or equal to 6

OpenStudy (anonymous):

sqrt (12) + 4x >= 6 4x >= 6 sqrt(12) x >= 6/4 sqrt(12)

OpenStudy (anonymous):

its sqrt (12+4x) ≥ 6 @dauspex

OpenStudy (anonymous):

(12+4x) >= 6^2 12+4x >= 36 4x >= 24 x >= ?

OpenStudy (anonymous):

6 :)